What quantity x of a 75% acid solution must be mixed with a 25% solution to produce 500 mL of a 40% solution?


Answer by Alan3354(69443)

What quantity x of a 75% acid solution must be mixed with a 25% solution to produce 500 mL of a 40% solution?
---------------------------
s = 75% acid volume
t = 25% acid volume
--------------
s + t = 500
0.75s + 0.25t = 0.4*500 = 200 --> x4
3s + t = 800
1s + t = 500
-------------- Subtract
2s = 300
s = 150 ml of 75%
t = 350 ml of 25%